OpenStudy (idealist10):

Find a particular solution of y"+y=(-4+8x)cosx+(8-4x)sinx.

OpenStudy (idealist10):

\[y _{p}=x[(A+Bx)cosx+(C+Dx)sinx]\]

OpenStudy (idealist10):

Is my initial guess right above?

OpenStudy (solomonzelman):

The solution to the corresponding homogeneous equation is: \(\color{#0000ff}{ \displaystyle y_p(x)=c_1\cos(x) + c_2\sin( x) }\) So, if you were to just write: \(\color{#0000ff}{ \displaystyle y_p(x)=(A_1x+A_2)\cos(x) +(A_3x+A_4) \sin( x) }\) Then, you are repeating the homogeneous solution. \(\color{#0000ff}{ \displaystyle y_p(x)=A_1x\cos(x)+\color{red}{\underline{\color{blue}{A_2\cos(x)}}} +A_3x\sin( x)+\color{red}{\underline{\color{blue}{A_4 \sin( x)}}} }\) For this reason, (like you did, good job) you add another linear factor \(x\). So, your initial guess is, indeed like you said: \(\color{#0000ff}{ \displaystyle y_p(x)=\color{green}{\underline{\color{blue}{x}}}(A_1x+A_2)\cos(x) +\color{green}{\underline{\color{blue}{x}}}(A_3x+A_4) \sin( x) }\)

OpenStudy (solomonzelman):

I am using \(A_i\) for coefficients, to track them better, if there are a lot of them, I hope that isn't a problem:) Now, since this initial guess is a solution to the differential equation (except that you don't know the coefficients yet), you have to plug this initial guess into your differential equation, and find your coefficients.

OpenStudy (solomonzelman):

good luck!

OpenStudy (idealist10):

Thank you for the help, it worked!